OpenStudy (eric_d):

Why does the total energy of atom x decreases when there is a transfer of an electron from one subshell to another @aaronq

OpenStudy (abhisar):

Hello @eric_d !

OpenStudy (abhisar):

When an electron gets transferred from a lower subshell to upper subshell then its kinetic energy decreases but its total energy increases

OpenStudy (abhisar):

\[TE = -13.6*\frac{Z ^{2}}{n^{2}}ev/atom\]

OpenStudy (abhisar):

here, n is the number of subshell. Now u can see that greater will be the value for n lesser will be the value of total energy in negative which implies that more will be the value of total energy

OpenStudy (abhisar):

U didn't specified in ur question if it's upstepping or downstepping of electron ? So total energy will decrease in case of downstepping

OpenStudy (abhisar):

u getting my points ?

OpenStudy (eric_d):

Yes, there are two possibilities for this question! Upstepping & Downstepping

OpenStudy (eric_d):

Don't mind explaning this... ' which implies that more will be the value of total energy' @Abhisar

OpenStudy (abhisar):

is the question clear to uh now ?

OpenStudy (eric_d):

yes

OpenStudy (abhisar):

That's great ! \(\color{green}{\huge\ddot\smile}\)

OpenStudy (eric_d):

So, the answer is..due to increase in K.energy as the electrons are transferred frm higher subshell to a lower subshell Is it so

OpenStudy (abhisar):

Total energy decreases while stepping down

OpenStudy (abhisar):

do u want me to xplain why it is so ?

OpenStudy (eric_d):

yes

OpenStudy (abhisar):

ohkay, what part in my above xplanation is giving you trouble ?

OpenStudy (eric_d):

just want to clarify this first kinetic.e is opposite of T.Energy

OpenStudy (abhisar):

Yes the ratio of K.E : T.E for any electron is -1:1

OpenStudy (eric_d):

which implies that more will be the value of total energy ?

OpenStudy (abhisar):

Total energy is maximum for an electron when it's at n=\(\infty\), minimum when it is at n=1.

OpenStudy (abhisar):

Kinetic energy is maximum for an electron when it at n=1, minimum when it is at n=\(\infty\)

OpenStudy (abhisar):

got it ?

OpenStudy (eric_d):

So, the answer is.. kinetic energy increases

OpenStudy (abhisar):

answer for what ?

OpenStudy (abhisar):

you mean to say that total energy decreases because kinetic energy increases ?

OpenStudy (eric_d):

Yes

OpenStudy (abhisar):

yes u can say that

OpenStudy (abhisar):

Basically T.E = P.E + K.E, now P.E = -\(\frac{KZe^{2}}{r}\) and K.E = \(\frac{KZe^2}{2r}\)

OpenStudy (abhisar):

Now when u will add these up u will get the first formula i told uh

OpenStudy (abhisar):

and i explained how first formula says that T.E decreases with a decrease in shell number and increases with increase in shell number

OpenStudy (abhisar):

I hope it's not getting more confusing for uh !

OpenStudy (eric_d):

Thanks for the explanation. What is the other option besides increase in K.E for this question

OpenStudy (abhisar):

I am not getting ur question !!

OpenStudy (eric_d):

So, there is only one answer for this question

OpenStudy (eric_d):

which is increase in Kinetic.E

OpenStudy (abhisar):

See it's always best to explain these questions mathematically on the basis of a valid formula. But if u want to answer it conceptually u can say that Total energy increases because the sum of kinetic energy and potential energy increases.

OpenStudy (abhisar):

is it clear ?

OpenStudy (eric_d):

Yes, it's very clear!

OpenStudy (eric_d):

Thanks @Abhisar

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You are Most Welcomed! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (somy):

cute and colorful <3 @Abhisar good job!!!

OpenStudy (abhisar):

Thnx @Somy !